# Differentially Private Bilevel Optimization

### v2

@@ 8 Proofs
$\left\|\nabla F_{\mathcal{P}}(\mathbf{x})-\nabla F_{S}(\mathbf{x})\right\|=\widetilde{\mathcal{O}}\left(G\sqrt{{d_{x}\log(D/\gamma)}/{n}}\right)~{},$
where $G$ is the Lipschitz constant of $F$666 Mei et al. (2018) originally stated this for functions whose gradients are sub-Gaussian vectors. By Lemma 8.9, the gradients are $G$-bounded, hence $\mathcal{O}(G)$-sub-Gaussian.
@@
With probability at least $21-\gamma/$ 2 it holds that
@@
$=\left\|\mathcal{G}_{F_{S},\eta}(\mathbf{x}_{\mathrm{out}})\right\|+\widetilde{\mathcal{O}}\left(\ell\kappa\sqrt{\frac{d_{x}}{n}}\right)~{},$
where the second inequality is due to Lemma 2.4, and the last is by Lemma 8.10 with the Lipschitz bound of Lemma 8.9 and the domain bound $\|\mathbf{x}_{\mathrm{out}}-\mathbf{x}_{0}\|\leq D$ for some sufficiently large $D$ which is polynomial in all problem parameters (therefore only affecting log terms). The results then follow from Theorems 4.1 and 5.1. ∎
@@ 9 Discussion
Our work leaves open several directions for future research. First, the convergence rate derived in this work might be improvable. Specifically, for “single”-level DP nonconvex optimization, Arora et al. (2023) showed that variance reduction leads to faster convergence to private stationarity. Applying this for DP bilevel optimization as the outer loop would require, according to our analysis, to evaluate the cost of inexact gradients in variance-reduced methods, which we leave for future work.
Another direction for future work is extending the analysis in Section 7 to derive update rules for privately tuning multiple hyperparameters simultaneously. It is interesting to note that the complexity of our results scales polynomially with the upper-level dimension, which corresponds to the number of hyperparameters, whereas preforming a grid search scales exponentially with the number of hyperparameters.
@@ Acknowledgments
This research is supported through an Azrieli Foundation graduate fellowship. We would like to express our gratitude to Vitaly Feldman, Kunal Talwar and the Apple MLR team for the encouraging environment in which this research took place. Specifically, we thank Kunal Talwar for insightful discussions, which included pointing out the utility of localizing DP-SGD for optimal high probability guarantees.
Appendix A Optimal DP algorithm for strongly-convex objectives
@@
In particular, we analyze localized DP (S)GD. Although it would have been more natural to apply DP-SGD, this seems (at least according to our analysis) to yield an inferior rate with respect to the required high probability guarantee.777 Even if we would have sought only expectation bounds with respect to the hyperobjective, the high probability bound with respect to the inner problem is key to being able to argue about the gradient inexactness thereafter. Indeed, for DP-(S)GD, previous works (such as Bassily et al. 2014) typically provide bounds in expectation, and then convert them into high-probability bounds via a black-box reduction, which applies several runs and selects the best run via the private noisy-min (i.e. Laplace mechanism). The additional error incurred by this selection is of order $\frac{1}{n}$, which translates to $\frac{1}{\sqrt{n}}$ in terms of distance to the optimum, thus spoiling the fast rate of $\frac{1}{n}$ otherwise achieved in expectation for strongly-convex objectives. We therefore resort to localization [Feldman et al., 2020]: by running projected-(S)GD over balls with shrinking radii, applying martingale concentration bounds enables us to show that the distance to optimum shrinks as $R_{m+1}\lesssim\sqrt{\frac{R_{m}}{n}}+\frac{1}{n}$, and thus with negligible overhead we eventually recover the optimal fast rate $R_{M}\lesssim\frac{1}{n}$ with high probability. Our analysis differs than previous localization analyses, as it does not require adapting the noise-level and step sizes throughout the rounds.
@@
Proof of Theorem A.1.
We start by proving the privacy guarantee. By the Lipschitz assumption, the sensitivity of $\nabla h(\cdot;B_{t})$ is at most $\frac{2L}{b}$, thus the “unamplified” Gaussian mechanism (Lemma 2.2) ensures $(\tilde{\epsilon},\tilde{\delta})$-DP with $\tilde{\epsilon}=\widetilde{\Theta}\left(\frac{L}{b\sigma}\right)=\widetilde{\Theta}\left(\frac{\epsilon^{\prime}}{b}\right)$, and hence is amplified (Lemma 2.3) to $(\epsilon_{0},\delta_{0})$-DP for $\epsilon_{0}=\widetilde{\Theta}\left(\frac{\epsilon^{\prime}}{b}\cdot\frac{b}{n}\right)=\widetilde{\Theta}\left(\frac{\epsilon^{\prime}}{n}\right)=\widetilde{\Theta}\left(\frac{\epsilon^{\prime}}{\sqrt{T}}\right)$. Advanced composition (Lemma 2.1) therefore ensures that the overall algorithm is $(\epsilon^{\prime},\delta^{\prime})$-DP (note that this uses the fact that $M=\widetilde{\mathcal{O}}(1)$).
@@
Appendix B Auxiliary Lemma
See 2.4
Proof of Lemma 2.4.

### v3

@@ 8 Proofs
[…]
$+\frac{(\ell^{3/2}\kappa^{5/2}+\ell^{1/2}\kappa^{3/2})\log^{1/4}(1/\gamma)}{n^{1/4}}\bigg)~.$
Proof of Lemma 8.9.
We start by setting up notation. Given a dataset $S\sim\mathcal{P}^{n}$, recall that we denote by $F_{\mathcal{P}}/F_{S}$ the stochastic/empirical hyperobjectives, by $f_{\mathcal{P}}/f_{S}$ and $g_{\mathcal{P}}/g_{S}$ the stochastic/empirical upper- and lower-level objectives. We denote by $\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x})=\arg\min_{\mathbf{y}}g_{\mathcal{P}}(\mathbf{x},\mathbf{y})$ and $\mathbf{y}^{*}_{S}(\mathbf{x})=\arg\min_{\mathbf{y}}g_{S}(\mathbf{x},\mathbf{y})$. It holds that
$\nabla F_{\mathcal{P}}(\mathbf{x})$ $=\nabla_{x}f_{\mathcal{P}}(\mathbf{x},\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x}))+(\frac{d\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x})}{d\mathbf{x}})^{\top}\nabla_{y}f_{\mathcal{P}}(\mathbf{x},\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x}))$
$\nabla F_{S}(\mathbf{x})$ $=\nabla_{x}f_{S}(\mathbf{x},\mathbf{y}^{*}_{S}(\mathbf{x}))+(\frac{d\mathbf{y}^{*}_{S}(\mathbf{x})}{d\mathbf{x}})^{\top}\nabla_{y}f_{S}(\mathbf{x},\mathbf{y}^{*}_{S}(\mathbf{x}))~.$
Thus,
$\|\nabla F_{\mathcal{P}}(\mathbf{x})-\nabla F_{S}(\mathbf{x})\|$ $\leq\|\nabla_{x}f_{\mathcal{P}}(\mathbf{x},\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x}))-\nabla_{x}f_{S}(\mathbf{x},\mathbf{y}^{*}_{S}(\mathbf{x}))\|$
$~~~+\|(\frac{d\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x})}{d\mathbf{x}})^{\top}\nabla_{y}f_{\mathcal{P}}(\mathbf{x},\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x}))-(\frac{d\mathbf{y}^{*}_{S}(\mathbf{x})}{d\mathbf{x}})^{\top}\nabla_{y}f_{S}(\mathbf{x},\mathbf{y}^{*}_{S}(\mathbf{x}))\|$
$\leq\underset{(1)}{\underbrace{\|\nabla_{x}f_{\mathcal{P}}(\mathbf{x},\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x}))-\nabla_{x}f_{S}(\mathbf{x},\mathbf{y}^{*}_{S}(\mathbf{x}))\|}}$
$~~~+\underset{(2)}{\underbrace{\|\frac{d\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x})}{d\mathbf{x}}\|\cdot\|\nabla_{y}f_{\mathcal{P}}(\mathbf{x},\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x}))-\nabla_{y}f_{S}(\mathbf{x},\mathbf{y}^{*}_{S}(\mathbf{x}))\|}}$
$~~~+\underset{(3)}{\underbrace{\|\nabla_{y}f_{S}(\mathbf{x},\mathbf{y}^{*}_{S}(\mathbf{x}))\|\cdot\|\frac{d\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x})}{d\mathbf{x}}-\frac{d\mathbf{y}^{*}_{S}(\mathbf{x})}{d\mathbf{x}}\|}}~$,
where the last inequality used the fact that $|ab-cd|=|ab-ad+ad-cd|\leq|a||b-d|+|d||a-c|$. We turn to bound each summand with high probability, which immediately results in Lemma 8.9 by summing and union bounding.
Bound $(1)$.
By smoothness, it holds that
$\|\nabla_{x}f_{\mathcal{P}}(\mathbf{x},\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x}))-\nabla_{x}f_{S}(\mathbf{x},\mathbf{y}^{*}_{S}(\mathbf{x}))\|\leq\|\nabla_{x}f_{\mathcal{P}}(\mathbf{x},\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x}))-\nabla_{x}f_{S}(\mathbf{x},\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x}))\|+L_{1}^{f}\|\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x})-\mathbf{y}^{*}_{S}(\mathbf{x})\|~.$
To bound the first term, we note that the partial derivatives are bounded by $L_{0}^{f}$ by Lipschitzness. Hence, we can apply a uniform convergence bound due to Mei et al. (2018, Theorem 1) and get with probability at least $1-\gamma/9:$
@@
$\|\nabla_{x}f_{\mathcal{P}}(\mathbf{x},\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x}))-\nabla_{x}f_{S}(\mathbf{x},\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x}))\|=\widetilde{\mathcal{O}}\left(L_{0}^{f}\sqrt{{d_{x}\log(D/\gamma)}/{n}}\right)~.$
To bound the second term, we use strong-convexity to get
$L_{1}^{f}\|\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x})-\mathbf{y}^{*}_{S}(\mathbf{x})\|\leq\frac{L_{1}^{f}}{\sqrt{\mu_{g}}}\sqrt{g_{\mathcal{P}}(\mathbf{x},\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x}))-g_{\mathcal{P}}(\mathbf{x},\mathbf{y}^{*}_{S}(\mathbf{x}))}~.$
Now, we note that since $\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x}),\mathbf{y}^{*}_{S}(\mathbf{x})$ are the stochastic/empirical minima of the strongly-convex stochastic optimization problem given by $g_{\mathcal{P}}$, a result due to Feldman and Vondrak (2019) bounds the generalization gap with probability at least $1-\gamma/9$ by
$g_{\mathcal{P}}(\mathbf{x},\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x}))-g_{\mathcal{P}}(\mathbf{x},\mathbf{y}^{*}_{S}(\mathbf{x}))=\widetilde{\mathcal{O}}\left(\frac{(L_{0}^{g})^{2}\log(1/\gamma)}{\mu_{g}n}+\frac{\sqrt{\log(1/\gamma)}}{\sqrt{n}}\right)~.$
Overall, by union bounding and simplifying the condition dependent constants, we get
$(1)=\widetilde{\mathcal{O}}\left(\frac{\ell\sqrt{{d_{x}\log(D/\gamma)}}}{\sqrt{n}}+\frac{\ell\kappa\sqrt{\log(1/\gamma)}}{\sqrt{n}}+\frac{\sqrt{\ell\kappa}\log^{1/4}(1/\gamma)}{n^{1/4}}\right)~.$
Bound $(2)$.
On one hand, by the implicit function theorem,
$\|\frac{d\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x})}{d\mathbf{x}}\|=\|\nabla^{2}_{xy}g_{\mathcal{P}}(\mathbf{x},\mathbf{y}_{\mathcal{P}}^{*}(\mathbf{x}))[\nabla^{2}_{yy}g_{\mathcal{P}}(\mathbf{x},\mathbf{y}_{\mathcal{P}}^{*}(\mathbf{x}))]^{-1}\|\leq L_{1}^{g}/\mu_{g}\leq\kappa~.$
Moreover, repeating the argument used to bound $(1)$, we get
$\|\nabla_{y}f_{\mathcal{P}}(\mathbf{x},\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x}))-\nabla_{y}f_{S}(\mathbf{x},\mathbf{y}^{*}_{S}(\mathbf{x}))\|$ $\leq\|\nabla_{y}f_{\mathcal{P}}(\mathbf{x},\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x}))-\nabla_{y}f_{S}(\mathbf{x},\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x}))\|+L_{1}^{f}\|\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x})-\mathbf{y}^{*}_{S}(\mathbf{x})\|$
$=\widetilde{\mathcal{O}}\left(\frac{\ell\sqrt{{d_{x}\log(D/\gamma)}}}{\sqrt{n}}+\frac{\ell\kappa\sqrt{\log(1/\gamma)}}{\sqrt{n}}+\frac{\sqrt{\ell\kappa}\log^{1/4}(1/\gamma)}{n^{1/4}}\right)~.$
Overall, we get
$(2)=\widetilde{\mathcal{O}}\left(\frac{\ell\kappa\sqrt{{d_{x}\log(D/\gamma)}}}{\sqrt{n}}+\frac{\ell\kappa^{2}\sqrt{\log(1/\gamma)}}{\sqrt{n}}+\frac{\ell^{1/2}{\kappa}^{3/2}\log^{1/4}(1/\gamma)}{n^{1/4}}\right)~.$
Bound $(3)$.
We first note that $\|\nabla_{y}f_{S}(\mathbf{x},\mathbf{y}^{*}_{S}(\mathbf{x}))\|\leq L_{0}^{f}\leq\ell$. Moreover, due to the implicit function theorem, the basic estimate $|ab-cd|\leq|a||b-d|+|d||a-c|$, the fact $\|A^{-1}-B^{-1}\|\leq\frac{1}{\lambda_{\min}(A)\lambda_{\min}(B)}\|A-B\|$, and second-order smoothness of $g$, we get
$\|\frac{d\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x})}{d\mathbf{x}}-\frac{d\mathbf{y}^{*}_{S}(\mathbf{x})}{d\mathbf{x}}\|$ $=\|\nabla^{2}_{xy}g_{\mathcal{P}}(\mathbf{x},\mathbf{y}_{\mathcal{P}}^{*}(\mathbf{x}))[\nabla^{2}_{yy}g_{\mathcal{P}}(\mathbf{x},\mathbf{y}_{\mathcal{P}}^{*}(\mathbf{x}))]^{-1}-\nabla^{2}_{xy}g_{S}(\mathbf{x},\mathbf{y}_{S}^{*}(\mathbf{x}))[\nabla^{2}_{yy}g_{S}(\mathbf{x},\mathbf{y}_{S}^{*}(\mathbf{x}))]^{-1}\|$
$\leq\|\nabla^{2}_{xy}g_{\mathcal{P}}(\mathbf{x},\mathbf{y}_{\mathcal{P}}^{*}(\mathbf{x}))\|\cdot\|[\nabla^{2}_{yy}g_{\mathcal{P}}(\mathbf{x},\mathbf{y}_{\mathcal{P}}^{*}(\mathbf{x}))]^{-1}-[\nabla^{2}_{yy}g_{S}(\mathbf{x},\mathbf{y}_{S}^{*}(\mathbf{x}))]^{-1}\|$
$~~~~+\|[\nabla^{2}_{yy}g_{S}(\mathbf{x},\mathbf{y}_{S}^{*}(\mathbf{x}))]^{-1}\|\cdot\|\nabla^{2}_{xy}g_{\mathcal{P}}(\mathbf{x},\mathbf{y}_{\mathcal{P}}^{*}(\mathbf{x}))-\nabla^{2}_{xy}g_{S}(\mathbf{x},\mathbf{y}_{S}^{*}(\mathbf{x}))\|$
$\leq(\frac{L_{1}^{g}}{\mu_{g}^{2}}+\frac{1}{\mu_{g}})\|\nabla^{2}g_{\mathcal{P}}(\mathbf{x},\mathbf{y}_{\mathcal{P}}^{*}(\mathbf{x}))-\nabla^{2}g_{S}(\mathbf{x},\mathbf{y}_{S}^{*}(\mathbf{x}))\|$
$\leq(\frac{L_{1}^{g}}{\mu_{g}^{2}}+\frac{1}{\mu_{g}})(\|\nabla^{2}g_{\mathcal{P}}(\mathbf{x},\mathbf{y}_{\mathcal{P}}^{*}(\mathbf{x}))-\nabla^{2}g_{S}(\mathbf{x},\mathbf{y}_{\mathcal{P}}^{*}(\mathbf{x}))\|+L_{2}^{g}\|\mathbf{y}_{\mathcal{P}}^{*}(\mathbf{x})-\mathbf{y}_{S}^{*}(\mathbf{x})\|)~.$
Applying a Hessian uniform convergence bound due to Mei et al. (2018, Theorem 1.b) gives with probability at least $1-\gamma/9:$
$\|\nabla^{2}g_{\mathcal{P}}(\mathbf{x},\mathbf{y}_{\mathcal{P}}^{*}(\mathbf{x}))-\nabla^{2}g_{S}(\mathbf{x},\mathbf{y}_{\mathcal{P}}^{*}(\mathbf{x}))\|=\widetilde{\mathcal{O}}\left(L_{1}^{g}\sqrt{{d_{x}\log(D/\gamma)}/{n}}\right)~.$
The same argument that appeared in bounding $(1)$ and $(2)$ gives
$L_{2}^{g}\|\mathbf{y}_{\mathcal{P}}^{*}(\mathbf{x})-\mathbf{y}_{S}^{*}(\mathbf{x})\|=\widetilde{\mathcal{O}}\left(\frac{\ell\kappa\sqrt{\log(1/\gamma)}}{\sqrt{n}}+\frac{\sqrt{\ell\kappa}\log^{1/4}(1/\gamma)}{n^{1/4}}\right)~.$
Combining all the pieces and simplifying the condition dependent constants, we get overall
$(3)=\widetilde{\mathcal{O}}\left(\frac{(\ell^{2}\kappa^{2}+\ell\kappa)\sqrt{d_{x}\log(D/\gamma)}}{\sqrt{n}}+\frac{(\ell^{2}\kappa^{3}+\ell\kappa^{2})\sqrt{\log(1/\gamma)}}{\sqrt{n}}+\frac{(\ell^{3/2}\kappa^{5/2}+\ell^{1/2}\kappa^{3/2})\log^{1/4}(1/\gamma)}{n^{1/4}}\right)~.$
With Lemma 8.9 in hand, we turn to prove Theorem 6.1.
@@
Using the fact that the gradient mapping is non-expansive (Lemma C.1), we see that with probability at least $1-\gamma/2$ the following holds:
@@
$=\left\|\mathcal{G}_{F_{S},\eta}(\mathbf{x}_{\mathrm{out}})\right\|+\widetilde{\mathcal{O}}\bigg(\frac{(\ell^{2}\kappa^{2}+\ell\kappa)\sqrt{d_{x}\log(1/\gamma)}}{\sqrt{n}}+\frac{(\ell^{2}\kappa^{3}+\ell\kappa^{2})\sqrt{\log(1/\gamma)}}{\sqrt{n}}$
$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~+\frac{(\ell^{3/2}\kappa^{5/2}+\ell^{1/2}\kappa^{3/2})\log^{1/4}(1/\gamma)}{n^{1/4}}\bigg)~,$
where the alst inequality is by Lemma 8.9 with a domain bound $\|\mathbf{x}_{\mathrm{out}}-\mathbf{x}_{0}\|\leq D$ for some sufficiently large $D$ which is polynomial in all problem parameters (therefore only affecting log terms). Theorem 6.1 follows from applying Theorems 4.1 and 5.1 to bound $\left\|\mathcal{G}_{F_{S},\eta}(\mathbf{x}_{\mathrm{out}})\right\|$ and simplifying.
@@ 9 Discussion
Our work leaves open several directions for future research. First, it is natural to ask what is the extent to which our results can be improved. Notably, even in the well-studied setting of single-level smooth-nonconvex DP optimization, there still exist gaps between known upper and lower bounds for minimizing the gradient norm (cf. Lowy et al. 2024 and discussion therein). The best known lower bound for such problems, which trivially applies also for DP BO which is a strictly more general problem setting, is $\Omega(\sqrt{d}/\epsilon n)$, hinting that our results are perhaps not tight. Moreover, Lowy and Liu (2025) recently presented improved gradient bounds for DP BO via second-order methods, and it is interesting to ask whether first-order methods for this setting, as studied in this work, can be improved.
To follow up on this question, we remark that a candidate strategy to improve the convergence rate in this work would be to use variance reduction, as Arora et al. (2023) used variance reduction to derive faster convergence to private stationarity for single-level DP nonconvex optimization. Applying this for DP BO as the outer loop would require, according to our analysis, to evaluate the cost of inexact gradients in variance-reduced methods, which is left for future work.
Another direction for future work is extending the analysis in Section 7 to derive update rules for privately tuning multiple hyperparameters simultaneously. It is interesting to note that the complexity of our results scales polynomially with the upper-level dimension, which corresponds to the number of hyperparameters, whereas preforming a grid search scales exponentially with the number of hyperparameters. The trade-off is, of course, convergence to local stationarity instead of global optimality.
@@ Acknowledgments
This research is supported through an Azrieli Foundation graduate fellowship. We would like to express our gratitude to Vitaly Feldman, Kunal Talwar and the Apple MLR team for the encouraging environment in which this research was initiated. Specifically, we thank Kunal Talwar for insightful discussions, which included pointing out the utility of localizing DP-SGD for optimal high probability guarantees.
Appendix A Differential privacy preliminaries
We recall here some well-known DP facts. The basic composition property of DP states that the (possibly adaptive) composition of $(\epsilon_{0},\delta_{0})$-DP- and $(\epsilon_{1},\delta_{1})$-DP mechanisms, is $(\epsilon_{0}+\epsilon_{1},\delta_{0}+\delta_{1})$-DP. Advanced composition improves this scaling in typical parameter regimes of interest:
Lemma A.1 (Advanced composition, Dwork et al., 2010).
We remark that advanced composition is typically used when $\epsilon_{0}\lesssim\sqrt{\log(1/\delta_{0})/T}$, thus the accumulated privacy scales as $\epsilon\asymp\sqrt{T}\epsilon_{0}$.
Lemma A.2 (Gaussian mechanism).
Lemma A.3 (Privacy amplification, Balle et al., 2018).
Privacy amplification is typically used when $\epsilon_{0}\leq 1$, under which the privacy after subsampling scales as $\epsilon\asymp\frac{b\epsilon_{0}}{n}$ (since $e^{\epsilon_{0}}-1\asymp\epsilon_{0},~(1-1/n)^{b}\asymp\frac{b}{n}$ and $\log(1+\frac{b}{n}\epsilon_{0})\asymp\frac{b}{n}\epsilon_{0}$).
Appendix B Optimal DP algorithm for strongly-convex objectives
@@
In particular, we analyze localized DP (S)GD. Although it would have been more natural to apply DP-SGD, this seems (at least according to our analysis) to yield an inferior rate with respect to the required high probability guarantee. [^7] Even if we would have sought only expectation bounds with respect to the hyperobjective, the high probability bound with respect to the inner problem is key to being able to argue about the gradient inexactness thereafter. Indeed, for DP-(S)GD, previous works (such as Bassily et al. 2014) typically provide bounds in expectation, and then convert them into high-probability bounds via a black-box reduction, which applies several runs and selects the best run via the private noisy-min (via Laplace mechanism). The additional error incurred by this selection is of order $\frac{1}{n}$, which translates to $\frac{1}{\sqrt{n}}$ in terms of distance to the optimum, thus spoiling the fast rate of $\frac{1}{n}$ otherwise achieved in expectation for strongly-convex objectives. We therefore resort to localization [Feldman et al., 2020]: by running projected-(S)GD over balls with shrinking radii, applying martingale concentration bounds enables us to show that the distance to optimum shrinks as $R_{m+1}\lesssim\sqrt{\frac{R_{m}}{n}}+\frac{1}{n}$, and thus with negligible overhead we eventually recover the optimal fast rate $R_{M}\lesssim\frac{1}{n}$ with high probability. Our analysis differs than previous localization analyses, as it does not require adapting the noise-level and step sizes throughout the rounds.
@@
Proof of Theorem B.1.
[…]
@@
Appendix C Auxiliary lemma
We will recall a useful fact, which asserts that the mapping $\mathcal{G}_{\mathbf{v},\eta}(\mathbf{x})$ is non-expansive with respect to $\mathbf{v}:$
Lemma C.1.
For any $\mathbf{x},\mathbf{v},\mathbf{w}\in\mathbb{R}^{d},~\eta>0:~\left\|\mathcal{G}_{\mathbf{v},\eta}(\mathbf{x})-\mathcal{G}_{\mathbf{w},\eta}(\mathbf{x})\right\|\leq\left\|\mathbf{v}-\mathbf{w}\right\|$.
In our analysis we argue that in the bilevel setting, gradient estimates must be inexact to avoid privacy leaking between the two levels, and Lemma C.1 allows us to control the error due to this inexactness. Lemma C.1 is known (cf. Ghadimi et al. 2016), and we reprove it here for completeness.
Proof of Lemma C.1.